# Towards the Usage of Window Counting Constraints

### v2

@@ Introduction
The automated translation of a system specification into its implementation is one of the most challenging problems in formal methods. Providing correct-by-construction artifacts, such synthesis offers great potential in the development of new systems by significantly reducing the need for manual work in the engineering process. In this article, we focus on synthesis for reactive systems, i.e. systems that are influenced by and interact with their environment. This interaction can be modeled as a game, in which the system tries to play according to its specification, whereas the moves of the environment can potentially impede the system from reaching its goal. Since the interaction between system and environment is typically of long-lasting nature without predefined end date, the game is infinite in the sense that a play of the game has infinite duration, while the arena, modeled as a graph, has finitely many states. The players play by moving a token from one state of the arena to the next. The player whose turn it is decides which of the outgoing transitions of the current state is chosen. A well-known type of game is the safety game: The system wins a play if it can avoid to reach predefined unsafe states. Otherwise, the environment wins. A player has a winning strategy, if it wins against all possible behavior of the other player. For two-player safety games on finite graphs, there always exists a winning strategy for one of the players and this winning strategy can be computed [BL69], [Tho95]. However, the efficient computation of winning strategies (not only in the case of safety games) is still an open challenge in the synthesis of reactive systems. A common synthesis approach is to first generate a deterministic word automaton as game graph from specifications written as Linear Temporal Logic (LTL) formulae. Then, a strategy that is winning in the game is calculated. By construction, the strategy automatically satisfies the specification. Unfortunately, the construction of the deterministic word automaton leads to an automaton with a number of states that is doubly exponential in the length of the specification [PR89], making the whole strategy synthesis unfeasible for many applications. For avoiding the most expensive part of the synthesis procedure, there exist synthesis algorithms that start with a subset of the specification language LTL, such that it is possible to construct the game graph in a more efficient way. One example for that is the usage of the LTL subclass Generalized Reactivity(1) (GR(1)), which allows to construct and solve the game in time $O(N^{3})$ with $N$ being the size of the state space [PPS06]. While GR(1) is expressive enough for the specification of many systems [MR15], some specifications that do not fall into GR(1) remain unconsidered. For example, Maoz and Ringert mention the consideration of synthesis with counting patterns as future work in [MR15], but to the best of our knowledge, this has not been addressed yet.
@@
In this article, we deal with the request for efficient synthesis for some types of counting patterns as part of the system specification and present the idea of recursive synthesis for such games. We call the considered counting patterns “window counting constraints”. These are of the form
@@
with parameters $k,l\in\mathbb{N},\,k\leq l$. The “at least” can also be replaced by “at most”. Such constraints arise naturally when the desired behavior of systems includes reoccurring elements. For example, an automated guided vehicle on a factory floor might need to charge its battery in at least two out of ten moves to avoid facing depleted batteries. By replacing “system” in the window counting constraint by “environment”, the constraint introduces rationality of the environment, since the environment will not primary behave antagonistic to the system in plays of the game, but will behave according to its own constraints. From the synthesis point of view, rationality of the environment implies information for the system on the environment, reducing the space of situations the system could face in a play of the game. The term “window” in the constraint-type name emphasizes the relation of those specifications to sliding windows in data stream monitoring [PS06]. For the sake of better readability, we also call them counting constraints for short.
@@
We avoid the direct full translation of the specifications into a graph and instead focus on the following two observations: (1) It is possible to influence how hard it is to satisfy a counting constraint by varying the parameters $k$, $l$ in the counting constraints. More precisely, the (non-)existence of a strategy that fulfills the specification in a game with a set of counting constraints allows to make statements about the (non-)existence of such a strategy in a game with a set of counting constraints with varied parameters. (2) The values in the counting constraints influence the scale of the game graph that encodes all information given by the constraints. The larger $k$ and $l$, the exponentially larger is the graph. Consequently, the values influence how much computational power and/or memory is needed in order to synthesize a winning strategy.
@@
Combining these observations, the presented approach can be summarized as follows: Consider a two-player game graph and some specifications in the form of counting constraints. For solving the synthesis problem of finding a strategy for the system, such that the counting constraints are fulfilled, start with a subset of counting constraints that result in a small game graph or a trivially winnable game. Calculate winning strategies (if existent) and check what the (non-)existence of a winning strategy means for a game with refined/relaxed (depending on the constraints) constraints. This information gives reliable hints on which parts of the game with adapted values in the counting constraints are worth to investigate in the next iteration step and which parts of the game graph can then be neglected, leading to a reduction in the state space. In each iteration, the set of considered counting constraints converges more to the game of interest. Although the size of the game graphs may increase in each iteration, the gained state space reduction leads to a synthesis algorithm more efficient than when considering the game of interest as a whole from the beginning. If the original game without window counting constraints is a reachability, safety, Büchi, co-Büchi or parity-game, then the synthesis task in each iteration can be done by any synthesis algorithm fitting to the game class. Our motivation for starting with a game graph accompanied with counting constraints instead of a pure set of specifications comes from the robotic domain. In many applications, automated guided vehicles are moving in specified areas (like a factory floor). Modeling the setting as a game graph in which states encode the position of systems arises naturally.
@@
This work is an extension of [FF24]. The extended version complements the content of the original one in two main areas: For once, the synthesis approach in [FF24] only considered safety games, while this work extends the game types to reachability, Büchi, co-Büchi and parity games. Secondly, this work adds the possibility of iterating over constraints of both the form “The system plays action actact at least $k$ times out of $l$ of its own moves.” and “The system plays action actact at most $k$ times out of $l$ of its own moves.” in the same synthesis procedure, while the previous version of the approach considered one of the constraint types to be fixed. Additionally, this article is enriched with more extensive explanations and examples.
@@
This article is structured as follows. In Section 1, related work in the field of synthesis for reactive systems is presented, focusing on the challenge of constructing efficient algorithms. After summarizing concepts and notations required to formulate the game, Section 2 provides the definition of a game with counting constraints. In Section 3, we present the idea of and an algorithm for iterated synthesis with counting constraints. An example and runtime results of the algorithm based on a non-optimized implementation are given in Section 4. Current limitations and chances of the iterated synthesis approach that guide our future work are collected in Section 5. Section 6 concludes the article.
@@ 1. Related Work
In 1957, Church formulated the Synthesis Problem as finding finite-memory procedures to transform an infinite sequence of input data into an infinite sequence of output data, such that the relation between input and output satisfies given specifications [Chu57], [Tho09]. Around a decade later, Büchi and Landweber showed the decidability of the problem [BL69]. However, the algorithmic complexity of synthesis algorithms remains a challenge. The translation of specifications from monadic second-order logic of one successor (S1S) into a Büchi automaton as part of the synthesis procedure is nonelementary in the length of specifications [Sto74]. This indicates that it is not possible to construct a universally (or an in all cases) efficient synthesis algorithm that can handle complete S1S specifications. For specifications expressible in Linear Temporal Logic (LTL), the problem is 2EXPTIME-complete [PR89].
@@
Acknowledging the absence of a generally low-complexity synthesis algorithm for arbitrary S1S/LTL specifications, the literature presents three primary approaches [Fin16]. The first approach restricts the scope of considered specifications for synthesis to less expressive logics. Here, the structure of the considered specifications is used to reduce the synthesis complexity. The second one is tackling the internal representation of the problem. Solutions following this approach are often aiming for algorithms with in average good runtime. In this approach, it suffices if most systems can be synthesized with acceptable resources (memory, computational time), while the existence of corner cases with worst-case complexity is accepted. The third approach focuses on the output of the problem, the implementation. The size of the implementation gets restricted such that only small implementations are accepted as solutions of the synthesis problem. The rationale behind this is that small and hence less complicated implementations often exist for applications. Such solutions are often easier (that is, with less computational time) identifiable than bigger (complex) implementations, if it is possible to steer the algorithm towards small solutions. Synthesis algorithms can follow more than one of those approaches.
@@
A well-studied class of specifications for approach (1) is General Reactivity of Rank 1 (GR1), a fragment of LTL for which there are symbolic synthesis algorithms that are polynomial in the size of the state space of the design [PPS06]. Examples for other specification classes for which efficient solutions of the synthesis problem are investigated are Safety LTL [ZTL+17], Metric Temporal Logic with a Bounded Horizon [MNP07] and Extended Bounded Response LTL [CGG+20].
@@
Following approach (2), Kupferman and Vardi developed a synthesis method that does not require the costly determinization of non-deterministic Büchi automata representing the specification [KV05], which is the most complex part in many synthesis algorithms. Other synthesis algorithms rely for instance on symbolic synthesis to represent sets of states of a game graph in a compact matter via antichains [FJR09], [FJR11], binary decision diagrams [Ehl10] and LTL fragements [CGG+20].
@@
The work by Schewe and Finkbeiner presents a synthesis algorithm that employs bounded synthesis as approach (3). Their method uses translation of LTL specifications into sequences of safety tree automata, in order to constrain the size of the implementation [SF07]. “Lazy synthesis”, in which an SMT solver is used to construct potential implementations for an incomplete constraint system, extends the system only if required [FJ12].
@@
The general idea is inspired by the work of Chen et al. on games with delay. In this work, one player only receives information on the moves of the environment with a delay of $k\in\mathbb{N}$ turns. For strategy synthesis, the delay is incrementally enlarged from zero to $k$ with a graph reduction step after each iteration step [CFL+21].
@@ 2. Games with Counting Constraints
[Two-player game graph] A two-player finite-state game graph (or: arena) is of the form $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$ where $S$ is a finite (non-empty) set of states, $S_{EGO}$, $S_{ALTER}$ define a partition of $S$, $s_{0}\in S_{EGO}$ is the initial state, $\Sigma_{EGO}$ is a finite alphabet of actions for player $EGO$, $\Sigma_{ALTER}$ is a finite alphabet of actions for player $ALTER$ and $\rightarrow\subseteq S\times(\mathcal{P}(\Sigma_{EGO})\cup\mathcal{P}(\Sigma_{ALTER}))\times S$ is a set of labeled transitions satisfying the following four conditions, where $\mathcal{P}(M)$ denotes the power set of a set $M$:
  • Bipartition: For each $(s,\sigma,s^{\prime})\in\rightarrow$ holds either (1) $s\in S_{EGO}$ and $s^{\prime}\in S_{ALTER}$ or (2) $s\in S_{ALTER}$ and $s^{\prime}\in S_{EGO}$.
@@
  • Alphabet restriction on actions: For a player $p\in\{EGO,ALTER\}$ holds: If $(s,\sigma,s^{\prime})\in\rightarrow$ with $s\in S_{p}$, then $\sigma\in\mathcal{P}(\Sigma_{p})$.
  • Determinacy of moves: For $p\in\{EGO,ALTER\}$ and $\sigma\in\mathcal{P}(\Sigma_{EGO})\cup\mathcal{P}(\Sigma_{ALTER})$ holds: if $s\in S_{p}$ and $(s,\sigma,s^{\prime}),(s,\sigma,s^{\prime\prime})\in\rightarrow$, then $s^{\prime}=s^{\prime\prime}$.
Such a game graph, also referred to as “arena”, encodes a game between the two players $EGO$ and $ALTER$. For $p\in\{EGO,ALTER\}$ the set of states $S_{p}$ contains the states where it is the turn of player $p$ to perform an action, also called $p$ controls $s$. Due to the bipartition and alphabetic restriction on actions, the game is “turn-based”, i.e. the two players alternate between choosing one of the possible actions. Since the game graph does not contain deadlocks, it results in an infinite sequence of states and actions, called an infinite play.
@@
[Infinite play] Let $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$ be a two-player game graph. An infinite play on $A$ is an infinite sequence $\pi=(\pi_{i}\sigma_{i})_{i\in\mathbb{N}_{0}}=\pi_{0}\sigma_{0}\pi_{1}\sigma_{1}\dots$ with $\pi_{0}=s_{0}$ and $\pi_{i}\sigma_{i}\pi_{i+1}\in\rightarrow$ for all $i\in\mathbb{N}_{0}$. $\Pi(A)$ denotes the set of all infinite plays on $A$.
If the played letters are not of relevance in a play, we also denote plays solely by the visited situations, i.e., $\pi\in S^{\omega}$. In such an infinite play, the two players play against (or in case of collaborative games: with) each other. Players can have strategies that determine how they react in each step of the play.
@@
  • For a play $\pi=(\pi_{i}\sigma_{i})_{i\in\mathbb{N}_{0}}$, a prefix of $\pi$ up to position $n$ is denoted by $\pi(n)=\pi_{0}\sigma_{0}\pi_{1}\dots\pi_{n-1}\sigma_{n-1}\pi_{n}$. The length of $\pi(n)$, denoted by $|\pi(n)|$, is $n+1$. The last state $\pi_{n}$ of $\pi(n)$ is called the tail of the prefix $\pi(n)$, denoted by $Tail(\pi(n))$. The set of all prefixes of plays in the game graph $A$ is $Pref(A)$.
  • For a player $p\in\{EGO,ALTER\}$ and a game graph $A$, the set of all prefixes that end in a state controlled by $p$ is $Pref_{p}(A):=\{\pi(n)\in Pref(A)\,|\,Tail(\pi(n))\in S_{p}\}$.
  • A strategy for a player $p\in\{EGO,ALTER\}$ in the game graph $A$ is a mapping $\varphi:\,Pref_{p}(A)\longrightarrow\mathcal{P}(\Sigma_{p})$, such that for all prefixes $\pi(n)\in Pref_{p}(A)$ and all $\sigma\in\varphi(\pi(n))$ there exist a state $s\in S\setminus S_{p}$ and a transition $(Tail(\pi(n)),\sigma,s)\in\rightarrow$.
  • The outcome $O(A,\varphi)$ of a strategy $\varphi$ of $p\in\{EGO,ALTER\}$ in the game graph $A$ is the set of all possible plays when player $p$ follows the strategy $\varphi$ and the other player plays arbitrary, i.e. $O(A,\varphi):=\{\pi=(\pi_{i}\sigma_{i})_{i\in\mathbb{N}_{0}}\in\Pi(A)\,|\,\forall i\in\mathbb{N}_{0}:\sigma_{2i}\in\varphi(\pi(2i))\text{ if }s_{0}\in S_{p}\text{ and }\sigma_{2i+1}\in\varphi(\pi(2i+1))\text{ otherwise}\}$.
A game $G=(A,\text{Win})$ consists of a game graph $A$ and a winning condition Win like a reachability, safety, Büch, co-Büchi or parity condition. The winning condition consist of a set of (infinite) sequences of states of $G$. EGOEGO wins if it has a strategy that guarantees to only plays in Win. The environment, on the other hand, wins if a play not in Win is played. Hence, each play of a two-player safety game always has exactly one winner and one loser. Games with this property are called zero sum games.
@@
[Winning conditions] Let $A$ be a game graph with states $S$ and $R\subseteq S$. $G=(A,\text{Win})$ becomes a game with any of the following winning conditions:
@@
  • The reachability winning condition is defined as $\text{Win}=\mathtt{Reach}(R):=\{\pi\in S^{\omega}\,|\,Occ(\pi)\cup R\neq\emptyset\}$. In a reachability game, player $EGO$ wins a play, if any state $s\in R$ is visited at least once in the play.
@@
We are now introducing window counting constraints as a mean to encode reoccurring behavior of a player with limits on which action can be selected how often in each snippet (or: window) of a play of a given length. Each window counting constraint belongs to a player (either $EGO$ or $ALTER$) and the action is a formula over the alphabet of the player. {defi}[Window Counting Constraints] Let $A$ be a game graph with the two players $EGO$ and $ALTER$, denote with $a$ propositional logic formula over the alphabet $\Sigma_{p}$ for $p\in\{EGO,ALTER\}$ and $k,l\in\mathbb{N}$ with $k\leq l$. For $\sigma\in\mathcal{P}(\Sigma_{p})$ we write $\sigma\models a$, if $\land_{b\in\sigma}b\models a$. Let $\pi=(\pi_{i}\sigma_{i})_{i\in\mathbb{N}_{0}}$ be a play on $G$.
@@
    For $p=EGO$, $CC_{max}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\leq k$.

### v3

@@ Introduction
The automated translation of a system specification into its implementation is one of the most challenging problems in formal methods. Providing correct-by-construction artifacts, such synthesis offers great potential in the development of new systems by significantly reducing the need for manual work in the engineering process. In this article, we focus on synthesis for reactive systems, i.e. systems that are influenced by and interact with their environment. This interaction can be modeled as a game, in which the system tries to play according to its specification, whereas the moves of the environment can potentially impede the system from reaching its goal. Since the interaction between system and environment is typically of long-lasting nature without predefined end date, the game is infinite in the sense that a play of the game has infinite duration, while the arena, modeled as a graph, has finitely many states. The players play by moving a token from one state of the arena to the next. The player whose turn it is decides which of the outgoing transitions of the current state is chosen. A well-known type of game is the safety game: The system wins a play if it can avoid to reach predefined unsafe states. Otherwise, the environment wins. A player has a winning strategy, if it wins against all possible behavior of the other player. For two-player safety games on finite graphs, there always exists a winning strategy for one of the players and this winning strategy can be computed [2], [24]. However, the efficient computation of winning strategies (not only in the case of safety games) is still an open challenge in the synthesis of reactive systems. A common synthesis approach is to first generate a deterministic word automaton as game graph from specifications written as Linear Temporal Logic (LTL) formulae. Then, a strategy that is winning in the game is calculated. By construction, the strategy automatically satisfies the specification. Unfortunately, the construction of the deterministic word automaton leads to an automaton with a number of states that is doubly exponential in the length of the specification [20], making the whole strategy synthesis unfeasible for many applications. For avoiding the most expensive part of the synthesis procedure, there exist synthesis algorithms that start with a subset of the specification language LTL, such that it is possible to construct the game graph in a more efficient way. One example for that is the usage of the LTL subclass Generalized Reactivity(1) (GR(1)), which allows to construct and solve the game in time $O(N^{3})$ with $N$ being the size of the state space [19]. While GR(1) is expressive enough for the specification of many systems [17], some specifications that do not fall into GR(1) remain unconsidered. For example, Maoz and Ringert mention the consideration of synthesis with counting patterns as future work in [17], but to the best of our knowledge, this has not been addressed yet.
@@
In this article, we deal with the request for heuristically efficient synthesis for some types of counting patterns as part of the system specification and present the idea of recursive synthesis for such games. The translation of the full system specification into a deterministic automaton as game arena is out of scope. We already start with a deterministic automaton accompanied with a set of counting patterns. The automaton can be obtained from a specification by any translation procedure (without encoding the counting patterns as well) or can arise naturally like it often occurs in the robotics domain, where the graph encodes the position of an automated system in specified areas like a factory floor. We call the considered counting patterns “window counting constraints”. These are of the form
@@
with parameters $k,l\in\mathbb{N},\,k\leq l$ and actact encoding an action controllable by the system. The “at least” can also be replaced by “at most”. Such constraints arise naturally when the desired behavior of systems includes reoccurring elements. For example, an automated guided vehicle on a factory floor might need to charge its battery in at least two out of ten moves to avoid facing depleted batteries. By replacing “system” in the window counting constraint by “environment”, the constraint introduces rationality of the environment, since the environment will not primary behave antagonistically to the system in plays of the game, but will behave according to its own constraints. From the synthesis point of view, rationality of the environment implies information for the system on the environment, reducing the space of situations the system could face in a play of the game. The term “window” in the constraint-type name emphasizes the relation of those specifications to sliding windows in data stream monitoring [18]. For the sake of readability, we also call them counting constraints for short.
@@
We avoid the direct full translation of the specifications in form of counting constraints into a graph and instead focus on the following two observations: (1) It is possible to influence how hard it is to satisfy a counting constraint by varying the parameters $k$, $l$ in the counting constraints. More precisely, the (non-)existence of a strategy that fulfills the specification in a game with a set of counting constraints allows to make statements about the (non-)existence of such a strategy in a game with a set of counting constraints with varied parameters. (2) The values in the counting constraints influence the scale of the game graph that encodes all information given by the constraints. The larger $k$ and $l$, the exponentially larger in the (unary) values of $k$ and $l$ is the graph, if the memory on the last moves is encoded into the states of the graph. Consequently, the values influence how much computational power and/or memory is needed in order to synthesize a winning strategy.
@@
Combining these observations, the presented approach can be summarized as follows: Consider a two-player game graph and some specifications in the form of counting constraints. For solving the synthesis problem of finding a strategy for the system, such that the counting constraints are fulfilled, start with a subset of counting constraints that result in a small game graph or a trivially winnable game. Calculate winning strategies (if existent) and check what the (non-)existence of a winning strategy means for a game with refined/relaxed (depending on the constraints) constraints. This information gives reliable hints on which parts of the game with adapted values in the counting constraints are worth to investigate in the next iteration step and which parts of the game graph can then be neglected, leading to a reduction in the state space. With each iteration, the set of considered counting constraints converges to the game of interest. Although the size of the game graphs may increase in each iteration, the gained state space reduction leads to a synthesis algorithm heuristically more efficient than when considering the game of interest as a whole from the beginning. If the original game without window counting constraints is a reachability, safety, Büchi, co-Büchi or parity-game, then the synthesis task in each iteration can be done by any synthesis algorithm fitting to the game class. This approach is not improving the complexity class of the synthesis problem, but shall serve as suggestion on how to improve existing synthesis algorithms heuristically for games that use counting constraints.
@@
This work is an extension of [9]. The extended version complements the content of the original one in two main areas: First, the synthesis approach in [9] only considered safety games, while this work extends the game types to reachability, Büchi, co-Büchi and parity games. Second, this work adds the possibility of iterating over constraints of both the form “The system plays action actact at least $k$ times out of $l$ of its own moves.” and “The system plays action actact at most $k$ times out of $l$ of its own moves.” in the same synthesis procedure, while the previous version of the approach considered one of the constraint types to be fixed. Additionally, this article is enriched with more extensive explanations and examples.
@@
This article is structured as follows. In Section 1, related work in the field of synthesis for reactive systems is presented, focusing on the challenge of constructing efficient algorithms. After summarizing concepts and notations required to formulate the game, Section 2 provides the definition of a game with counting constraints. In Section 3, we present the idea of incremental synthesis with counting constraints and an algorithm for it. An example and runtime results of the algorithm based on a non-optimized implementation are given in Section 4. Current limitations and chances of the incremental synthesis approach that guide our future work are collected in Section 5. Section 6 concludes the article.
@@ 1. Related Work
In 1957, Church formulated the Synthesis Problem as finding finite-memory procedures to transform an infinite sequence of input data into an infinite sequence of output data, such that the relation between input and output satisfies given specifications [5], [25]. Around a decade later, Büchi and Landweber showed the decidability of the problem [2]. However, the algorithmic complexity of synthesis algorithms remains a challenge. The translation of specifications from monadic second-order logic of one successor (S1S) into a Büchi automaton as part of the synthesis procedure is nonelementary in the length of specifications [23]. This indicates that it is not possible to construct a universally efficient synthesis algorithm that can handle complete S1S specifications. For specifications expressible in Linear Temporal Logic (LTL), the problem is 2EXPTIME-complete [20].
@@
Acknowledging the absence of a generally low-complexity synthesis algorithm for arbitrary S1S/LTL specifications, the literature presents three primary approaches [13]. The first approach restricts the scope of considered specifications for synthesis to less expressive logics. Here, the structure of the considered specifications is used to reduce the synthesis complexity. The second one is tackling the internal representation of the problem. Solutions following this approach are often aiming for algorithms with in average good runtime. In this approach, it suffices if most systems can be synthesized with acceptable resources (memory, computational time), while the existence of corner cases with worst-case complexity is accepted. The third approach focuses on the output of the problem, the implementation. The size of the implementation gets restricted such that only small implementations are accepted as solutions of the synthesis problem. The rationale behind this is that small and hence less complicated implementations often exist for applications. Such solutions are often easier (that is, with less computational time) identifiable than bigger (complex) implementations, if it is possible to steer the algorithm towards small solutions. Synthesis algorithms can follow more than one of those approaches.
@@
A well-studied class of specifications for approach (1) is General Reactivity of Rank 1 (GR1), a fragment of LTL for which there are symbolic synthesis algorithms that are polynomial in the size of the state space of the design [19]. Examples for other specification classes for which efficient solutions of the synthesis problem are investigated are Safety LTL [26], Metric Temporal Logic with a Bounded Horizon [16] and Extended Bounded Response LTL [6].
@@
Following approach (2), Kupferman and Vardi developed a synthesis method that does not require the costly determinization of non-deterministic Büchi automata representing the specification [14], which is the most complex part in many synthesis algorithms. Other synthesis algorithms rely for instance on symbolic synthesis to represent sets of states of a game graph in a compact matter via antichains [10], [11], binary decision diagrams [8] and LTL fragments [6].
@@
The work by Schewe and Finkbeiner presents a synthesis algorithm that employs bounded synthesis as approach (3). Their method uses translation of LTL specifications into sequences of safety tree automata, in order to constrain the size of the implementation [22]. “Lazy synthesis”, in which an SMT solver is used to construct potential implementations for an incomplete constraint system, extends the system only if required [12].
@@
The general idea is inspired by the work of Chen et al. on safety games with delay. In this work, one player only receives information on the moves of the environment with a delay of $k\in\mathbb{N}$ turns. In particular, the player is forced to decide on its next move without having knowledge on the latest k turns of the recent history of the play. The key idea is that a strategy that is not winning for the player for some delay $l\leq k$ will also not be winning in the game with delay $k$. Moreover, moves of the system that are not part of a winning strategy for the game with delay $l$ will also not be part of a winning strategy for the game with larger delay. Hence, such behavior does not need to be considered for larger delays, allowing to reduce the required transition system that encodes the game with a larger delay (pruning step). The synthesis algorithm begins with assuming a delay of zero (no delay), which is the well-known problem of solving a safety game. If a winning strategy for the player exists, the constructed transition system is pruned and the delay is increased by one for the next iteration. If no winning strategy exists in any iteration for delay $l\leq k$, no winning strategy for delay $k$ exists [4].
@@ 2. Games with Counting Constraints
[Two-player game graph] A two-player finite-state game graph (or: arena) is of the form $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$ where $S$ is a finite (non-empty) set of states, $S_{EGO}$, $S_{ALTER}$ define a partition of $S$, $s_{0}\in S_{EGO}$ is the initial state, $\Sigma_{EGO}$ is a finite alphabet of actions for player $EGO$, $\Sigma_{ALTER}$ is a finite alphabet of actions for player $ALTER$ and $\rightarrow\subseteq\bigcup_{p\in\{EGO,ALTER}\}S_{p}\times\mathcal{P}(\Sigma_{p})\times S_{\overline{p}}$ is a set of labeled transitions satisfying the following two conditions, where $\mathcal{P}(M)$ denotes the power set of a set $M$, $\overline{EGO}=ALTER$ and $\overline{ALTER}=EGO$:
@@
  • Determinism of moves: For all $(s,\sigma,s^{\prime}),(s,\sigma,s^{\prime\prime})\in\rightarrow$ holds $s^{\prime}=s^{\prime\prime}$.
Such a game graph, also referred to as “arena”, encodes a game between the two players $EGO$ and $ALTER$. For $p\in\{EGO,ALTER\}$ the set of states $S_{p}$ contains the states where it is the turn of player $p$ to perform an action, also called $p$ controls $s$. The game is “turn-based”, i.e. the two players alternate between choosing one of the possible actions. Since the game graph does not contain deadlocks, it results in an infinite sequence of states and actions, called an infinite play.
@@
[Infinite play, prefix] Let $A=(S,s_{0},S_{EGO},S_{ALTER},\Sigma_{EGO},\Sigma_{ALTER},\rightarrow)$ be a two-player game graph. An infinite play on $A$ is an infinite sequence $\pi=(\pi_{i}\sigma_{i})_{i\in\mathbb{N}_{0}}=\pi_{0}\sigma_{0}\pi_{1}\sigma_{1}\dots$ with $\pi_{0}=s_{0}$ and $\pi_{i}\sigma_{i}\pi_{i+1}\in\rightarrow$ for all $i\in\mathbb{N}_{0}$. $\Pi(A)$ denotes the set of all infinite plays on $A$. A finite prefix of an infinite play $\pi$ is $h=(\pi_{i}\sigma_{i})_{0\leq i\leq n}=\pi_{0}\sigma_{0}\pi_{1}\sigma_{1}\dots\sigma_{n-1}\pi_{n}$ (for some $n\in\mathbb{N}_{0}$). It ends in $Tail(h)=\pi_{n}$, called its tail. It belongs to $EGO$ if $\pi_{n}\in S_{EGO}$ and to $ALTER$ otherwise. The set of prefixes of $EGO$ (respectively, $ALTER$) is denoted $\mathrm{Pref}_{EGO}(A)$ (resp., $\mathrm{Pref}_{ALTER}(A)$), and the set of prefixes is $\mathrm{Pref}=\mathrm{Pref}_{EGO}\uplus\mathrm{Pref}_{ALTER}$.
Since the two players alternate in choosing their actions, we can also interpret a play $\pi\in\Pi(A)$ as interleaving $\pi=u\otimes v:=(u_{i}v_{i})_{i\in\mathbb{N}_{0}}$ of unique sequences $u=(u_{i})_{i\in\mathbb{N}_{0}}\in(S_{EGO}\times\mathcal{P}(\Sigma_{EGO}))^{\omega}$, $v=(v_{i})_{i\in\mathbb{N}_{0}}\in(S_{ALTER}\times\mathcal{P}(\Sigma_{ALTER}))^{\omega}$. A prefix can we interpreted likewise. If the played letters are not of relevance in a play, we also denote plays solely by the visited situations, i.e., $\pi\in S^{\omega}$. In such an infinite play, the two players play against (or in case of collaborative games: with) each other. Players can have strategies that determine how they react in each step of the play.
@@
  • A strategy for a player $p\in\{EGO,ALTER\}$ in the game graph $A$ is a mapping $\varphi:\,\mathrm{Pref}_{p}(A)\longrightarrow\mathcal{P}(\Sigma_{p})$, such that for all prefixes $h\in\mathrm{Pref}_{p}(A)$ and all $\sigma\in\varphi(h)$ there exist a state $s\in S\setminus S_{p}$ and a transition $(Tail(h),\sigma,s)\in\rightarrow$.
  • A play $\pi$ on a game graph $A$ is an outcome of a strategy $\varphi$ of $p\in\{EGO,ALTER\}$ in the game graph $A$ when player $p$ follows the strategy $\varphi$ and the other player plays arbitrarily, i.e.: for all prefixes $h\pi_{n}\sigma_{n}$ (for some $\pi_{n}\in S$ and $\sigma_{n}\in\mathcal{P}(\Sigma)$) that are prefixes of $\pi$, if $h\in\mathrm{Pref}_{p}(A)$, then $\sigma_{n}=\varphi(h\pi_{n})$ (and if $h\in\mathrm{Pref}_{\overline{p}}(A)$, no condition is imposed). The set of outcomes of $\varphi$ on game graph $A$ is denoted $O(A,\varphi)$.
  • A strategy $\varphi$ for player $p\in\{EGO,ALTER\}$ in the game graph $A$ is positional, if the players next move only depends on the current state. In particular, a positional strategy requires no memory on the prefix of a run except for the current state. Formally: $\varphi$ is positional $\Leftrightarrow\forall\pi,\omega\in\mathrm{Pref}_{p}(A):Tail(\pi)=Tail(\omega)\Rightarrow\varphi(\pi)=\varphi(\omega)$.
A game $G=(A,\text{Win})$ consists of a game graph $A$ and a winning condition Win like a reachability, safety, Büchi, co-Büchi or parity condition. The winning condition consist of a set of (infinite) sequences of states of $G$. EGOEGO wins if it has a strategy $\varphi$ that guarantees $O(A,\varphi)\subseteq\text{Win}$. The environment, on the other hand, wins if a play not in Win is played. Hence, each play of such a two-player game with $\omega$-regular winning condition always has exactly one winner and one loser. Games with this property are called zero sum games.
In this article, safety games are used in all examples. Note that the presented approach is applicable to other winning conditions like the ones below. Required winning condition-specific modifications are described in the explanation of the approach (Section 3).
@@
[Winning conditions] Let $A$ be a game graph with states $S$ and $R\subseteq S$. For a play $\pi$ on $A$ denote the set of states of $A$ that are visited at least once in $\pi$ with $Occ(\pi)$ and the set of states that are visited infinitely often in $\pi$ as $Inf(\pi)$. $G=(A,\text{Win})$ becomes a game with any of the following winning conditions:
@@
  • The reachability winning condition is defined as $\text{Win}=\mathtt{Reach}(R):=\{\pi\in S^{\omega}\,|\,Occ(\pi)\cap R\neq\emptyset\}$. In a reachability game, player $EGO$ wins a play, if any state $s\in R$ is visited at least once in the play.
@@
A game with any of the above winning conditions is positionally determined, that is, for any state of the game, one of the players has a positional winning strategy.
We are now introducing window counting constraints as a mean to encode reoccurring behavior of a player with limits on which action can be selected how often in each snippet (or: window) of a play of a given length. Each window counting constraint belongs to a player (either $EGO$ or $ALTER$) and the action is a formula over the alphabet of the player.
{defi}
[Window Counting Constraints] Let $A$ be a game graph with the two players $EGO$ and $ALTER$. Let $p\in{EGO,ALTER}$. Let $\pi=u^{EGO}\otimes u^{ALTER}=(u^{EGO}_{i}u^{ALTER}_{i})_{i\in\mathbb{N}_{0}}$ be a play on $G$. Denote with $a$ a propositional logic formula over the alphabet $\Sigma_{p}$ and $k,l\in\mathbb{N}$ with $k\leq l$. Define for $i\in\mathbb{N}_{0}$ and $u^{p}_{i}=(s,\sigma)\in S_{p}\times\mathcal{P}(\Sigma_{p})$ $u^{p}_{i}\models a:\Leftrightarrow\land_{b\in\sigma}b\models a$.
@@
    $CC_{max}(p,a,k,l)$ is defined as the abbreviation for “The player $p$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(p,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{m\in\mathbb{N}_{0}\,|\,u^{p}_{m}\models a,\,i\leq m\leq i+l\}|\leq k$.